The sum of two numbers is 120. If one of the numbers is four times the other, find the smaller number.
Answer by ankor@dixie-net.com(22740)

The sum of two numbers is 120.
a + b = 120
If one of the numbers is four times the other,
a = 4b
find the smaller number.
:
In the 1st equation, replace a with 4b
4b + b = 120
5b = 120
b = 120/5
b = 24 is the smaller number
:
:
Check
a = 4(24) = 96
96 + 24 = 120